OpenStudy (anonymous):

y varies indirectly as twice x. When x = 2, y = 3. Find y when x = 6?!?! please help!!

OpenStudy (anonymous):

"Y varies indirectly as X" means that Y decreases as X increases and can be written as y=k/x where k is some constant.

OpenStudy (anonymous):

So when we read that "y varies indirectly as twice x," we can write y = k/(2x)

OpenStudy (anonymous):

Now, it's important to understand that k is some constant number. We just don't know what it is yet. If we did know what it was, then we could plug in any x, and find out what y was for that x.

OpenStudy (anonymous):

BUT, there's a way to solve for k, and once we solve for k, we can use the equation to solve for y when x=6.

OpenStudy (anonymous):

thank you!! :)

OpenStudy (anonymous):

My pleasure. Did you get it?

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

What was your answer, so I can check it?

OpenStudy (anonymous):

um well im not sure if its right but i got 2??

OpenStudy (anonymous):

Hmmm... What did you get for k?

OpenStudy (anonymous):

Your question should not be the way u asked, twice x should not be used as the 2 can go to the constant of proportionality

OpenStudy (anonymous):

I don't think there's anything wrong with the wording of her question. I think it just needs to be interpreted as y = k/(2x) for some k.

OpenStudy (anonymous):

ok fine then c=k/2